starla has been investing $5,000 quarterly for the past 12 years in a fixed income mutual fund. how much is the fund worth now assuming she has earned 4% compounded quarterly on her investment?
Answer #1
Answer: $306,113.04
What is required is the FV of the annuity of $5000 per quarter, deposited
for 12 years, that is 12*4 = 48 quarters, at 4% compounded quarterly.
FV = 5000*(1.01^48-1)/(0.01) = $   306,113.04